# Time-VLM: Exploring Multimodal Vision-Language Models for Augmented Time Series Forecasting

### v1

@@ 4.1 Few-shot Forecasting
We evaluate the few-shot capabilities of Time-VLM by testing its performance using only 5% or 10% of training data. This assesses its ability to combine pre-trained multimodal knowledge from VLM with time series-specific features for effective forecasting under minimal task-specific data.
As shown in Table 1 and Table 2, Time-VLM consistently outperforms most baselines across datasets. For example, on ETTh1 with 5% training data, Time-VLM reduces MSE by 29.5% and MAE by 16.6% compared to the second-best model, TimeLLM. On ETTm1 with 10% data, it surpasses TimeLLM by 11.1% in MSE and 10.5% in MAE. On Weather with 5% data, Time-VLM outperforms TimeLLM by 7.7% in MSE and 9.4% in MAE.
The performance gap between Time-VLM and traditional models (e.g., PatchTST, FEDformer) is more pronounced in few-shot settings, demonstrating the superiority of multi-modality in data-scarce scenarios. Notably, Time-VLM achieves this with only 143M parameters, significantly fewer than TimeLLM’s 3405M, highlighting its efficiency.
@@ 4.2 Zero-shot Forecasting
We evaluate the zero-shot capability of Time-VLM in cross-domain settings, where the model predicts on unseen datasets by effectively leveraging knowledge from unrelated domains. To ensure a more comprehensive and rigorous comparison, we use the ETT datasets as Time-LLM (Jin et al., 2024), with results summarized in Table 3.
Time-VLM demonstrates strong zero-shot generalization, consistently outperforming or matching state-of-the-art baselines with fewer parameters. For example, in $ETTh1\rightarrow ETTh2$, Time-VLM surpasses TimeLLM with a 4.2% lower MSE and 0.5% lower MAE. In $ETTm1\rightarrow ETTh2$, it outperforms TimeLLM by 7.1% in MSE and 3.6% in MAE. In $ETTm2\rightarrow ETTh2$, Time-VLM achieves competitive performance, closely matching TimeLLM with only a 1.4% difference in MSE and 0.3% in MAE.
@@ 4.3 Short-term Forecasting
For short-term forecasting, we evaluate Time-VLM on the M4 benchmark, which includes marketing data at various sampling frequencies. Performance is measured using SMAPE, MASE, and OWA metrics, averaged across datasets and sampling intervals (see Table 4).
Time-VLM demonstrates strong performance, consistently outperforming state-of-the-art baselines across all metrics. For instance, it surpasses the second-best model, Time-LLM, with improvements of 0.7% in SMAPE, 0.2% in MASE, and 0.5% in OWA, all while utilizing significantly fewer parameters and computational resources. Compared to traditional models like PatchTST and N-HiTS, the performance gains more, highlighting the benefit of multimodal knowledge in short-term forecasting. These gains stem from Time-VLM’s integration of temporal, visual, and textual data, capturing richer features for improved accuracy.
@@ 4.4 Long-term Forecasting
We evaluate the long-term forecasting capabilities of Time-VLM across diverse temporal horizons and datasets.
As shown in Table 5, Time-VLM achieves competitive performance compared to state-of-the-art baselines. For example, on ETTh1, Time-VLM surpasses TimeLLM with 0.7% improvements in MSE and MAE. On ETTm2, it outperforms TimeLLM by 1.2% in MSE and 0.6% in MAE. However, on Weather, Time-VLM slightly trails TimeLLM with a 0.4% higher MSE and 2.3% higher MAE.
@@ 4.5 Model Analysis
Ablation Studies: Table 6 evaluates the contributions of key components of Time-VLM, including the RAL, VAL, and TAL. Results are averaged across forecasting horizons $H\in\{96,192,336,720\}$ on the Weather dataset, with performance degradation (%Deg) measured for each variant.
Table 6: Ablation study on multimodal components.
Horizon Full w/o RAL w/o VAL w/o TAL
MSE MAE MSE MAE MSE MAE MSE MAE
96 0.160 0.213 0.273 0.324 0.213 0.266 0.165 0.218
192 0.203 0.252 0.297 0.338 0.237 0.298 0.208 0.257
336 0.253 0.291 0.325 0.354 0.255 0.302 0.258 0.295
720 0.317 0.340 0.369 0.383 0.309 0.357 0.322 0.345
Avg 0.233 0.274 0.316 0.350 0.254 0.306 0.238 0.279
%Deg $35.6\%\uparrow$ $27.7\%\uparrow$ $9.0\%\uparrow$ $11.7\%\uparrow$ $2.1\%\uparrow$ $1.8\%\uparrow$
@@
The study highlights the critical role of each component. Removing the RAL causes the largest performance drop ($35.6\%$ in MSE and $27.7\%$ in MAE), emphasizing its importance in capturing temporal dependencies through memory bank interactions. The VAL, which transforms time series into visual representations, is essential, with its exclusion leading to significant degradation ($9.0\%$ in MSE and $11.7\%$ in MAE). This underscores its ability to preserve fine-grained temporal patterns using VLM vision encoder. In contrast, removing the TAL results in minor degradation ($2.1\%$ in MSE and $1.8\%$ in MAE), likely due to sparse textual tokens in the VLM output (e.g., 11 out of 156 in ViLT). While the TAL provides valuable semantic context, its impact is limited by the VLM’s temporal understanding. Future work could explore larger VLMs with extended textual inputs to enhance temporal-semantic alignment.
@@
Multimodal and Few/Zero-shot Analysis: Time-VLM’s few-shot and zero-shot capabilities arise from its integration of temporal, visual, and textual modalities. The RAL models temporal dependencies through memory bank interactions, ensuring robust feature extraction with limited data. The VAL captures visually interpretable features (e.g., trend, seasonality, periodicity) in domain-agnostic visual representations, while the TAL generates contextual descriptions, providing semantic insights for better generalization. Together, these components enable Time-VLM to leverage pre-trained multimodal knowledge, making it highly adaptable to new tasks and domains with minimal training data.
@@
Figure 3: 2D UMAP visualization (Left) and Gate weight distributions (Right) of multimodal and temporal memory embeddings, highlighting their complementary behavior.
To validate the adaptation of VLM capabilities to time series, we analyze the similarity between RAL (temporal) and TAL/VAL (multimodal) embeddings. Figure 3 visualizes their complementary behavior. The left panel shows balanced gate weight distributions, indicating effective fusion of multimodal and temporal representations. The right panel’s UMAP visualization reveals distinct yet overlapping clusters, confirming successful integration of multimodal information while preserving unique characteristics. This demonstrates Time-VLM’s ability to adapt VLM-derived embeddings for robust time series analysis.
@@
Computation Studies: Time-VLM demonstrates strong computational efficiency, as shown in Table 7. With only 143.6M parameters (1/20 of Time-LLM’s 3404.6M), memory usage scales from 1968 MiB (Weather) to 24916 MiB (Traffic), adapting to dataset complexity. Inference speed ranges from 0.2057s/iter (ECL) to 0.4809s/iter (ETTh1), efficiently handling varying loads. In contrast, Time-LLM requires over 37GB of memory even for smaller datasets like ETTh1 and ETTh2, making it infeasible for larger datasets such as Weather, ECL, and Traffic. This highlights Time-VLM’s lightweight design and practical scalability.
@@
Hyperparameter Studies: We analyze key hyperparameters’ impact on performance, as shown in Figure 4. The sequence length performs best between 96 and 1024 timesteps, with 512 being optimal for most datasets. Longer sequences introduce noise without significant gains. The normalization constant peaks at 0.4, while the model dimension performs best at 128 for simpler datasets (e.g., ETTh1, ETTh2) and larger values for complex ones (e.g., Traffic, Weather). The gate network dimension, controlling multimodal fusion, achieves optimal results at 256 across most datasets.
@@ 5 Conclusion
We presented Time-VLM, a novel framework leveraging pretrained VLMs to unify temporal, visual, and textual modalities for time series forecasting. By integrating the RAL, VAL, and TAL, Time-VLM bridges modality gaps, enabling rich cross-modal interactions. Extensive experiments demonstrate state-of-the-art performance across various datasets, especially in few-shot and zero-shot scenarios, outperforming existing methods while maintaining efficiency. Our work establishes a new direction for multimodal time series forecasting, highlighting the potential of VLMs in capturing temporal dynamics and semantic context.
Notably, Time-VLM operates can solely on original time series data without external information, ensuring fair comparisons and showcasing its ability to generate textual and visual representations directly from the data for self-augmentation. This design not only enhances accuracy but also emphasizing the framework’s robustness, particularly in domains where external data is scarce or unavailable.
Future work may explore adaptive visual transformations for complex patterns, enhancing text utilization, extending to multi-task, and developing more efficient multimodal time series foundation models. For details, see Appendix D.
@@ C.2 Visualization of prediction results
The prediction results in Figures 6, 7, 8, and 9 demonstrate Time-VLM’s ability to accurately forecast time series across diverse datasets and prediction horizons. For datasets with clear periodic structures, such as the daily cycles in ETTh1 and ETTm1, Time-VLM captures both global trends and fine-grained temporal patterns effectively. This is evident in the close alignment between the true values (solid lines) and predicted values (dashed lines) across all horizons. Similarly, for the ECL dataset, which exhibits regular consumption patterns, Time-VLM delivers highly accurate forecasts, showcasing its strength in handling structured environments.

### v2

@@ 4.1 Few-shot Forecasting
Setting. We evaluate the few-shot long-term forecasting capabilities of Time-VLM by testing its performance using only 5% or 10% of the training data. This setting assesses how effectively Time-VLM integrates pre-trained multimodal knowledge from the VLM with time series-specific features under minimal task-specific supervision.
Results. As shown in Table 1 and Table 2, Time-VLM consistently outperforms most baselines across datasets. For example, on ETTh1 with 5% training data, Time-VLM reduces MSE by 29.5% and MAE by 16.6% compared to the second-best model, TimeLLM. On ETTm1 with 10% data, it surpasses TimeLLM by 11.1% in MSE and 10.5% in MAE. On Weather with 5% data, Time-VLM outperforms TimeLLM by 7.7% in MSE and 9.4% in MAE. The performance gap between Time-VLM and traditional models (e.g., PatchTST, FEDformer) is particularly pronounced in few-shot settings, demonstrating the effectiveness of multimodal integration when data is scarce. This performance gain stems from the model’s ability to leverage rich multimodal priors from pre-trained VLMs, while effectively capturing temporal patterns through memory-enhanced attention.
@@ 4.2 Zero-shot Forecasting
Setting. We evaluate the zero-shot forecasting capability of Time-VLM in cross-domain settings, where the model predicts on unseen datasets by effectively transferring knowledge from unrelated domains. To ensure a rigorous comparison, we conduct experiments using the ETT datasets as source and target domains, following previous setup (Jin et al., 2024). Results are summarized in Table 3.
Results. Time-VLM demonstrates strong generalizability, consistently outperforming or matching state-of-the-art baselines while using fewer parameters. For example, in the ETTh1->ETTh2 transfer setting, Time-VLM achieves a 4.2% lower MSE and 0.5% lower MAE than TimeLLM. In ETTm1->ETTh2, it outperforms TimeLLM by 7.1% in MSE and 3.6% in MAE. In ETTm2->ETTh2, Time-VLM performs competitively, closely matching TimeLLM with only a 1.4% difference in MSE and 0.3% in MAE. These results highlight Time-VLM’s ability to generalize across domains without fine-tuning, leveraging pre-trained vision-language priors for effective knowledge transfer.
@@ 4.3 Short-term Forecasting
Setting. For short-term forecasting, we evaluate Time-VLM on the M4 benchmark, which includes marketing data at various sampling frequencies. Performance is measured using SMAPE, MASE, and OWA metrics, averaged across datasets and sampling intervals (see Table 4).
Results. Time-VLM demonstrates strong performance, consistently outperforming state-of-the-art baselines across all metrics. For instance, it surpasses the second-best model, Time-LLM, with improvements of 0.7% in SMAPE, 0.2% in MASE, and 0.5% in OWA, all while utilizing significantly fewer parameters and computational resources. Compared to traditional models like PatchTST and N-HiTS, the performance gains more, highlighting the benefit of multimodal knowledge in short-term forecasting. These gains stem from Time-VLM’s integration of temporal, visual, and textual data, capturing richer features for improved accuracy.
@@ 4.4 Long-term Forecasting
Setting. We evaluate the long-term forecasting capabilities of Time-VLM across multiple horizons and datasets.
Results. As shown in Table 5, Time-VLM achieves competitive performance compared to state-of-the-art baselines. On ETTh1, Time-VLM improves upon TimeLLM by 0.7% in both MSE and MAE. On ETTm2, it outperforms TimeLLM by 1.2% in MSE and 0.6% in MAE. However, on the Weather dataset, Time-VLM slightly underperforms TimeLLM, with a 0.4% higher MSE and 2.3% higher MAE.
@@ 4.5 Model Analysis
Ablation Studies. Table 6 evaluates the contributions of key components of Time-VLM, including the RAL (with Local and Global Memory), VAL, and TAL. The study highlights the importance of each module. Removing the RAL causes a significant performance drop (35.6% in MSE), with its local (RAL_L) and global (RAL_G) branches contributing 17.2% and 4.3%, respectively, validating our hierarchical memory design. The VAL is also essential—removing it increases MSE by 9.0%, demonstrating its ability to preserve fine-grained temporal patterns via the VLM vision encoder. In contrast, removing the TAL results in only minor degradation (2.1% in MSE), likely due to the sparsity of textual tokens in the VLM output; for example, ViLT produces just 11 textual tokens out of 156 total, the rest being visual embeddings. While the TAL provides useful semantic context, its impact is limited by the scarcity of textual signals. Future work may explore VLMs with stronger language capabilities for better temporal-semantic alignment.
Table 6: Ablation study on multimodal components over forecasting horizons $H\in\{96,192,336,720\}$ on Weather dataset, with MSE performance degradation (%Deg) measured for each variant.
Horizon Full w/o RAL w/o RAL_L w/o RAL_G w/o VAL w/o TAL
96 0.160 0.273 0.185 0.165 0.213 0.165
192 0.203 0.297 0.235 0.210 0.237 0.208
336 0.253 0.325 0.295 0.265 0.255 0.258
720 0.317 0.369 0.375 0.330 0.309 0.322
Avg 0.233 0.316 0.273 0.243 0.254 0.238
%Deg $35.6\%\uparrow$ $17.2\%\uparrow$ $4.3\%\uparrow$ $9.0\%\uparrow$ $2.1\%\uparrow$
@@
Multimodal and Few-/Zero-shot Analysis. To understand the source of Time-VLM’s strong performance in data-scarce scenarios, we examine the relationship between RAL (temporal) and TAL/VAL (multimodal) embeddings. As shown in Figure 3, their complementary nature is evident. The left panel illustrates balanced gate weight distributions, indicating effective fusion of temporal and multimodal signals. The right panel presents a UMAP visualization revealing distinct yet partially overlapping clusters, confirming successful integration of multimodal information while preserving modality-specific characteristics. The robust few-shot and zero-shot capabilities of Time-VLM stem from its integration of temporal, visual, and textual modalities. Specifically, the RAL models temporal dependencies via memory bank interactions, enabling robust feature extraction even with limited data. The VAL captures interpretable visual patterns—such as trends, seasonality, and periodicity—in domain-agnostic representations, while the TAL generates semantic descriptions that enhance generalization. Together, these components allow Time-VLM to leverage pre-trained multimodal knowledge, making it highly adaptable to new tasks and domains with minimal fine-tuning.
@@
Figure 3: UMAP visualization (left) and gate weight distributions (right) of of multimodal and temporal embeddings.
Interpretability Analysis. We investigate how pre-trained VLMs can be leveraged for time series forecasting by analyzing the alignment between visual, textual, and temporal representations. To this end, we sample 400 image-text pairs from MSCOCO, the primary pre-training dataset for VLMs, and 200 samples each from time series datasets: ETT, Traffic, Weather, and ECL. Using UMAP, we visualize four types of embeddings in a shared 2D space:
  • multimodal embeddings derived from COCO-Pair samples using VLM, reflecting the model’s general pretrained cross-modal knowledge, which is Time-VLM’s main motivation of the time series project to.
  • multimodal embeddings from time series-generated image-text pairs processed through the same VLM, representing Time-VLM’s task-specific augmentation.
  • Visual-only embeddings from COCO-Image samples extracted via ViT, reflecting pure visual knowledge.
  • Text-only embeddings from COCO-Text samples encoded with BERT, capturing linguistic knowledge.
As shown in Figure 4, both COCO-Image and COCO-Text form distinct, separate clusters, remaining isolated from time-series features. This suggests that while low-level visual patterns in COCO-Image resemble temporal dynamics, pure image representations retain modality-specific characteristics. Similarly, COCO-Text forms a completely separate cluster, highlighting significant modality gaps. In contrast, COCO-Pair exhibits maximal overlap with time-series data, demonstrating strong cross-modal complementarity. The textual semantics in COCO-Pair bridge visual and temporal modalities, enhancing their alignment. Notably, COCO-Pair is positioned near the center of time-series clusters, suggesting its role as a key mediator between modalities.
These observations motivate Time-VLM’s design: instead of relying on single-modality projections, we embed time series into multimodal space for richer semantic understanding. The model’s strong performance in data-scarce settings stems from pre-trained VLM knowledge. With more time series data, better alignment of multimodal embeddings is expected, further enhancing performance.
Figure 4: Interpretability visualization of Time-VLM: multimodal feature alignment via UMAP.
@@
Computation Studies. Time-VLM demonstrates strong computational efficiency, as shown in Table 7. With only 143.6M parameters (1/20 of Time-LLM’s 3404.6M), memory usage scales from 1968 MiB (Weather) to 24916 MiB (Traffic), adapting to dataset complexity. Inference speed ranges from 0.2057s/iter (ECL) to 0.4809s/iter (ETTh1), efficiently handling varying loads. In contrast, Time-LLM requires over 37GB of memory even for smaller datasets like ETTh1 and ETTh2, making it infeasible for larger datasets such as Weather, ECL, and Traffic. This highlights Time-VLM’s lightweight design and practical scalability.
@@
Hyperparameter Studies. We analyze the impact of key hyperparameters on performance, as shown in Figure 5. Sequence length performs best between 96 and 1024 timesteps, with 512 being optimal for most datasets. Longer input introduce noise without significant gains, indicating that local temporal patterns are sufficient for accurate forecasting. The normalization constant peaks at 0.4, reflecting a balance between feature scaling and training stability. Model dimension shows dataset-dependent behavior: values of 128–256 suffice for short-term datasets like ETTh1 and ETTh2, while longer horizons and higher-dimensional data (e.g., Traffic, Weather) benefit from larger dimensions (up to 512), suggesting greater capacity is needed to model complex dynamics and variable interactions. Similarly, the gate network dimension—responsible for multimodal fusion—achieves optimal performance at 256 for medium-range forecasts. For more challenging settings like long-horizon or high-variable inputs, increasing it to 336 or 512 further improves results, highlighting the importance of adaptive fusion in capturing complex cross-modal relationships.
@@
VLM Variants Analysis. We conduct ablation studies on different VLM backbones and custom combinations to assess their impact on forecasting performance and computational efficiency. We evaluate three widely-used VLMs—ViLT, CLIP, and BLIP-2—with varying size, and find that increased model size does not improve accuracy. Notably, although BLIP-2 is the largest (3.7B parameters, 25GB+ memory), it underperforms ViLT and CLIP in terms of MSE (0.342 vs. 0.337) and exhibits slow training speed (0.98 s/iter), limiting its practical use. In contrast, lightweight models like ViLT (128.9M parameters, 1346 MiB memory) and CLIP (168.4M, 1174 MiB) achieve comparable or better accuracy at a fraction of the cost. To evaluate the benefit of VLMs’ pre-trained cross-modal alignment, we construct a modular baseline using separately trained vision and language encoders: ViT-B/16 and BERT-Base. As shown in Table 8, this custom combination underperforms all pretrained VLMs across metrics, achieving an average MSE of 0.348 versus 0.336 for ViLT, with no compensating gain in speed (0.17 s/iter). These findings highlight that the cross-modal alignment in VLMs offers a key inductive bias for time series modeling. Modular approaches, lacking such a unified space, fail to match this performance—demonstrating the value of pre-aligned multimodal representations for efficient fusion and forecasting.
Table 8: Comparison of different VLM variants on ETTh2 in terms of performance and computational efficiency.
VLM Type Params (M) Mem. (MiB) Speed (s/iter) MSE (avg) MAE (avg)
ViLT 128.9 1346 0.36 0.336 0.388
CLIP 168.4 1174 0.12 0.339 0.391
BLIP-2 3763.1 25200 0.98 0.342 0.393
Custom 213.2 1474 0.17 0.348 0.397
@@ 5 Conclusion
We presented Time-VLM, a novel framework that leverages pretrained VLMs to unify temporal, visual, and textual modalities for time series forecasting. By integrating the RAL, VAL, and TAL modules, Time-VLM bridges modality gaps and enables rich cross-modal interactions. Notably, it operates solely on raw time series data without requiring external information, enabling fair comparisons and demonstrating the ability to generate textual and visual representations internally for self-augmentation. This design not only improves accuracy but also highlights the framework’s robustness—particularly in domains where auxiliary data is scarce. Extensive experiments show that Time-VLM achieves superior performance across diverse datasets, especially in few-shot and zero-shot settings, outperforming existing methods while maintaining computational efficiency. Our work establishes a new direction in multimodal time series forecasting by highlighting the potential of VLMs in capturing both temporal dynamics and semantic context.
Limitations. Despite its strengths, Time-VLM has several limitations. First, the TAL module provides semantic context but has limited impact due to current VLMs’ constrained understanding of time series semantics. Second, while excelling in low-data regimes, full-shot performance slightly lags behind specialized unimodal models on certain tasks (e.g., ECL, Traffic), suggesting room for domain-specific adaptation. Third, although computationally efficient compared to LLM-based methods, deployment on resource-constrained devices remains challenging. These limitations suggest promising directions for future work, including temporally aware VLMs, improved time series imaging, visual distillation, and enhanced text encoders with stronger temporal reasoning. For details, see Appendix D.
Acknowledgements
This work is mainly supported by the Guangdong Basic and Applied Basic Research Foundation (No. 2025A1515011994). This work is also supported by the National Natural Science Foundation of China (No. 62402414, No. 62402420), Guangzhou Municipal Science and Technology Project (No. 2023A03J0011), the Guangzhou Industrial Information and Intelligent Key Laboratory Project (No. 2024A03J0628), and a grant from State Key Laboratory of Resources and Environmental Information System, and Guangdong Provincial Key Lab of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things (No. 2023B1212010007).
@@ C.2 Visualization of prediction results
The prediction results in Figures 7, 8, 9, and 10 demonstrate Time-VLM’s ability to accurately forecast time series across diverse datasets and prediction horizons. For datasets with clear periodic structures, such as the daily cycles in ETTh1 and ETTm1, Time-VLM captures both global trends and fine-grained temporal patterns effectively. This is evident in the close alignment between the true values (solid lines) and predicted values (dashed lines) across all horizons. Similarly, for the ECL dataset, which exhibits regular consumption patterns, Time-VLM delivers highly accurate forecasts, showcasing its strength in handling structured environments.